Tacoma's population in 2000 was about 200 thousand, and has been growing by about 9% each year. If this continues, what will Tacoma's population be in 2014?
Answer #1

Solution:

Given in the question

Population at the begining of 2000 was about p = 200000.

Rate of growth(r) = 9% each year

No. Of years(t) = 2014-2000 = 14 years

So growth equation

P0 = P*(1+(r/100))^t

P(14) = 200000*(1+(9/100))^14

P(14) = 200000*(1+0.09)^14

P(14) = 200000*(1.09)^14 = 200000*3.341727027 = 668345

So Tacoma population in 2014 will be 668345